OpenStudy (pounce0129):

Deanna and Pietra each go for walks around a lake a few times per week. Last week, Deanna walked 7 miles more than Pietra. a. If p represents the number of miles Pietra walked, write an equation that represents the total number of miles T the two girls walked. b. If Pietra walked 9 miles during the week, how many miles did Deanna walk? c. If Pietra walked 11 miles during the week, how many miles did the two girls walk together?

OpenStudy (anonymous):

a. T = p + (p +7 ) = 2p +7

OpenStudy (anonymous):

b . p = 9 deana = p+ 7 = 16

OpenStudy (anonymous):

c. p = 11 then T = 2p+7 = 22 + 7 = 29